# Local certification of geometric graph classes

### v4

@@ 6. Unit-square graphs
Let $G$ be a triangle-free unit-square graph embedded in the plane. For $n\geq 1$, a horizontal prop of length $n$ in $G$ is a sequence of distinct vertices $(u_{i})_{0\leq i\leq n}$, $(v_{i})_{0\leq i\leq n-1}$, and $(w_{i})_{0\leq i\leq n-1}$, such that the following holds: for each $10\leq i\leq n-$ 1, $v_{i}=n_{11}(u_{i})$, $w_{i}=n_{10}(u_{i})$, and $u_{i+1}=n_{10}(v_{i})=n_{11}(w_{i})$. Similarly, a vertical prop of length $n$ in $G$ is a sequence of distinct vertices $(u_{i})_{0\leq i\leq n}$, $(v_{i})_{0\leq i\leq n-1}$, and $(w_{i})_{0\leq i\leq n-1}$, such that the following holds: for each $10\leq i\leq n-$ 1, $v_{i}=n_{10}(u_{i})$, $w_{i}=n_{00}(u_{i})$, and $u_{i+1}=n_{00}(v_{i})=n_{10}(w_{i})$ (see Figure 6). The vertices $u_{0}$ and $u_{n}$ are respectively said to be the starting and ending vertex of the (horizontal or vertical) prop.
@@
Let $\mathrm{Pr}_{n}$ denote the graph induced by a (vertical or horizontal) prop of length $n$. That is, $\mathrm{Pr}_{n}$ can be obtained from $n$ disjoint 4-cycles $u_{i}v_{i}u_{i}^{\prime}w_{i}$ ($0\leq i\leq n-1)$ by identifying $u_{i}^{\prime}$ with $u_{i+1}$, for each $20\leq i\leq n-$ 2. Then $\mathrm{Pr}_{n}$ has $13n+$ 1 vertices. Consider any fixed embedding of $\mathrm{Pr}_{n}$ in the plane as a unit-square graph. Since $\mathrm{Pr}_{n}$ is triangle-free, this unit-square embedding of $\mathrm{Pr}_{n}$ also gives a planar embedding of $\mathrm{Pr}_{n}$ (with the same circular order of neighbors around each vertex), see Proposition 2.1. There are multiple non-equivalent planar embeddings of $\mathrm{Pr}_{n}$, however a simple area computation shows that in any planar graph drawing coming from a unit-square embedding of $\mathrm{Pr}_{n}$ each 4-cycle of $\mathrm{Pr}_{n}$ is a face, distinct from the outerface, so up to reflection the resulting planar embedding is unique. This implies the following.
@@
  • a cycle $C$ of length $1616k+$ 16, depicted with bold black edges in Figure 7;
@@
  • another copy of $\mathrm{Pr}_{8k+6}$ with endpoints $v_{2},v_{4}\in C$, such that $v_{2}$ and $v_{4}$ are antipodal on $C$ and, for any $1\leq i\leq 4$, $v_{i}$ and $v_{i+1}$ are at distance $44k+$ 4 on $C$ (where indices are taken modulo 4 plus 1). Moreover the two props intersect in their middle, forming a star on 5 vertices (depicted in bold red edges in Figure 7);
  • on the subpath of $C$ of length $44k+$ 4 between $v_{1}$ and $v_{2}$, there exists a unique vertex set $I$ of size $k$ in which all vertices are at distance at least 4 from $v_{1}$ and $v_{2}$ on $C$, and any two vertices of $I$ are at distance at least 4 on $C$. For each vertex $v\in I$, we create two copies of a 3 by 3 grid and add an edge between $v$ and a vertex of degree 3 in each of the two grids. The $2k$ copies of the grid are denoted by $H_{1},H_{1}^{\prime},\ldots,H_{k},H_{k}^{\prime}$ in order from $v_{1}$ and $v_{2}$.
@@
By Observation 6.3, $d_{\infty}(v_{1},v_{3})>8k+6$ and $d_{\infty}(v_{2},v_{4})>8k+6$. As $d_{\infty}(v_{i},v_{i+1})\leq 4k+4$ for any $1\leq i\leq 4$, this implies (by the triangle inequality) that $d_{\infty}(v_{i},v_{i+1})>4k+2$ for any such $i$ (with indices taken modulo 4 plus 1). Hence, we have proved that $44k+2<d_{\infty}(v_{1},v_{2})\leq 4k+$ 4. Using Observation 6.3, this implies that
@@
where $x(\cdot)$ and $y(\cdot)$ denote the $x$- and $y$-coordinates of the points as before. Let us denote by $u_{1},u_{2},\ldots,u_{4k+5}$ the vertices on the subpath $P$ of $C$ of length $44k+$ 4 between $v_{1}$ and $v_{2}$ (with $u_{1}=v_{1}$ and $u_{4k+5}=v_{2}$). The previous result implies that for any $51\leq i\leq j\leq 4k+$ 5,
@@
In particular, if we translate and rotate the vertices of the embedding of $G_{k}$ such that $v_{1}$ lies at coordinates $(0,0)$ and $v_{2}$ lies at $\ell_{\infty}$-distance at most 2 from $(4k+4,4k+4)$, then for any $51\leq i\leq 4k+$ 5, $u_{i}$ lies in the square with corners $(i-3,i-3)$ and $(i-1,i-1)$. We call this property the almost-perfect rigidity of $G_{k}$.
@@
The results obtained above show that for any $A,B\subseteq\{1,\ldots,N\}$, $g(L(A),R(B))$ is triangle-free, and it a unit-square graph if and only if $A$ and $B$ are disjoint. As $G_{N}^{\prime}$ and $G_{N}$ have $O(N)$ vertices and the closed neighborhood of $S$ is independent of $A$ and $B$ and has size 6, the class of triangle-free unit-square graphs is $(6,1)$-disjointness-expressing. ∎
@@ 7.3. Stripes
For $k,\delta>0$, a triangle in the plane is said to be $(k,\delta)$-almost-equilateral if all sides have length at least $k-\delta$ and at most $k+\delta$. By the law of cosines and the approximation $\arccos(1/2-x)=\pi/3+O(x)$ as $x\to 0$, we have the following.
@@
By Observation 7.1, the maximum angle between the lines $u_{i}v_{i}$ and $u_{i+1}v_{i+1}$ is of order $O(\delta/k)$. Hence, there exists a constant $\alpha$ (independent of $k$ and $\delta$) such that if $\ell=\alpha k/\delta$, the maximum angle between $u_{0}v_{0}$ and $u_{\ell}v_{\ell}$ in any $(k,\delta)$-almost-equilateral embedding is close to $\pi$. In this case, by Lemma 7.2 the minimum (Euclidean) distance $m$ between $u_{0}$ and $v_{\ell}$ in any $(k,\delta)$-almost-equilateral embedding is of order $\Theta(k/\delta\cdot k)=\Theta(k^{2}/\delta)$. Moreover, any $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$ realizing this minimum $m$ is close to some semicircle with endpoints $u_{0}$ and $v_{\ell}$, in the sense that all the vertices of $S_{\ell}$ lie at distance $O(k)$ from the semicircle (see Figure 11). We will need a looser version of this observation in the slightly weaker setting where $d_{2}(u_{0},v_{\ell})\leq m+1$, instead of $d_{2}(u_{0},v_{\ell})=m$.
@@
Let $k\gg\delta$ and $\ell=\lceil\alpha k/\delta\rceil$ as above, and let $m=\Theta(k^{2}/\delta)$ be the minimum Euclidean distance between $u_{0}$ and $v_{\ell}$ in any $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$. Consider a $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$ where $d_{2}(u_{0},v_{\ell})\leq m+1$. Let $c$ be the midpoint of the segment $[u_{0},v_{\ell}]$. Then no vertex of the stripe $S_{\ell}$ is contained in the disk of center $c$ and radius $m/2-O(k)$, and all vertices of the stripe $S_{\ell}$ are contained in a disk of center $c$ and radius $O(m)=O(k^{2}/\delta)$.
@@ 7.4. Quasi-rigid graphs
We define $T_{n}$ as follows. We consider two adjacent vertices $u,v$ of the infinite square grid and define $X$ as the set of vertices at distance at most $12n+$ 1 from $u$ or $v$ in the square grid, and $Y\subseteq X$ as the set of vertices at distance exactly $12n+$ 1 from $u$ or $v$. Note that $|X|=2(2n+2)^{2}=8(n+1)^{2}$ and $|Y|=8n+6$. The vertices of $Y$ are denoted by $y_{1},\ldots,y_{8n+6}$. The graph $T_{n}$ is obtained from the subgraph of the square grid induced by $X$ by adding, for each vertex $y_{i}$ of $Y\subseteq X$, a vertex $c_{i}$ adjacent to $y_{i}$. We finally add edges to form a cycle $C$ containing all vertices $c_{i}$ in order, together with 10 new vertices (2 or 3 at each corner, see Figure 12). The resulting cycle $C$ has length $168n+$ 16. Note that the resulting graph $T_{n}$ contains $8n^{2}+24n+24=O(n^{2})$ vertices and is a planar triangle-free unit-disk graph (see Figure 12). A simple area computation shows that in any unit-disk embedding of $T_{n}$, $C$ bounds the outerface of the corresponding planar graph embedding (by the arguments of Section 2.2 there is a unique planar graph embedding in the sphere, and as all faces except $C$ have size at most 8, $C$ must be the outerface).
@@
Let $f(n)$ be the infimum Euclidean distance between two antipodal vertices $c$ and $c^{\prime}$ of the cycle $C$ in a unit-disk embedding of $T_{n}$. By Lemma 7.4, $f(n)=\Omega(n)$. Since $C$ has length $O(n)$, it follows that $f(n)=\Theta(n)$. Assume for simplicity that the infimum $f(n)$ is a minimum (otherwise we work with a sequence of unit-disk embeddings such that the Euclidean distance between $c$ and $c^{\prime}$ tends to $f(n)$). Let $Z_{n}$ be the point set of a unit-disk embedding of $T_{n}$ in which $d_{2}(c,c^{\prime})=f(n)$. Add $\lceil f(n)\rceil-1$ points, evenly spaced on the line-segment $[c,c^{\prime}]$ (note that together with $c$ and $c^{\prime}$, any two consecutive points on $[c,c^{\prime}]$ lie at distance at most 1 apart). Let $Z_{n}^{\prime}$ denote the resulting point set and $T_{n}^{\prime}$ be the resulting unit-disk graph. Note that $T_{n}^{\prime}$ has $O(n^{2})$ vertices and in any unit-disk graph embedding of $T_{n}^{\prime}$, $f(n)\leq d_{2}(c,c^{\prime})\leq f(n)+1$, with $f(n)=\Omega(n)$.
@@
Assume we have found a family $\mathcal{T}^{1+c}$ which is quasi-rigid with density $g(k)=O(k^{1+c})$, for some $c\geq 0$. Take some graph $G\in\mathcal{T}^{1+c}$ with $g(k)$ vertices and that has two vertices $x$ and $y$ such that in any unit-disk embedding of $G$, $k\leq d_{2}(x,y)\leq k+1$. Let $X$ be the point set associated to some unit-disk embedding of $G$, and let $d:=d_{2}(x,y)$ in this point set ($k\leq d\leq k+1$). Consider a $(d,0)$-almost equilateral embedding of a stripe $S_{\ell}$ (for some integer $\ell\geq 0$ whose value will be fixed later), and for each edge $uv$ in $S_{\ell}$, add an isometric copy of $X$ in which $x$ is identified with $u$ and $y$ is identified with $v$ (by definition of $X$ we can always find such an isometric copy of $X$). See Figure 13 for an illustration. We denote by $X^{(\ell)}$ the resulting point set, and by $G^{(\ell)}$ the resulting unit-disk graph. Note that the different copies of $X$ interact and thus $G^{(\ell)}$ contains more edges than the unions of copies of $G$. Fix any unit-disk embedding of $G^{(\ell)}$, and consider the points corresponding to $x$ and $y$ in each of the copies of $G$ in $G^{(\ell)}$. By the properties of $G$, the union of all these points form a $(d,1)$-almost equilateral embedding of a stripe $S_{\ell}$. Let $a$ and $b$ be the ends of this stripe. By Lemma 7.3, there exists a constant $\alpha>0$ such that if we set $\ell=\lceil\alpha d\rceil=O(k)$, the following holds. Let $m$ be the minimum distance between $a$ and $b$ in any unit-disk embedding of $G^{(\ell)}$. Then $m=\Theta(d^{2})=\Theta(k^{2})$. As before we consider a point set realizing this distance and we add $\lceil m\rceil-1$ evenly spaced points on the segment $[a,b]$. We denote the resulting unit-disk graph by $H$ and we note that in any unit-disk graph embedding of $H$, $m\leq d_{2}(a,b)\leq m+1$. The point set $X^{(\ell)}$ has size at most $5\ell g(k)=O(k^{2+c})$ as there are $14\ell+$ 1 edges in $S_{\ell}$. We have added at most $m=O(d^{2})=O(k^{2})$ vertices to the graph, so $H$ has $O(k^{2+c})$ vertices. As $m=\Omega(k^{2})$, the family of all graphs $H$ created in this way is quasi-rigid with density $g^{\prime}(k)=O(k^{1+c/2})$.
@@
For any $\varepsilon>0$ there is a family of unit-disk graphs which is quasi-rigid with density $O(k^{1+\varepsilon})$. More precisely, for any sufficiently large $k$ there is a unit-disk graph $G$ with $O(k^{1+\varepsilon})$ vertices with two vertices $a$ and $b$ such that in any unit-disk embedding of $G$, $k\leq d_{2}(a,b)\leq k+1$, and $a$ and $b$ are joined by a path $P$ of length at most $k+1$ such that at least $k/2$ consecutive vertices of $P$ lie at Euclidean distance at least $\Omega(k)$ from $G-P$.

### v5

@@ 6. Unit-square graphs
Let $G$ be a triangle-free unit-square graph embedded in the plane. For $n\geq 1$, a horizontal prop of length $n$ in $G$ is a sequence of distinct vertices $(u_{i})_{0\leq i\leq n}$, $(v_{i})_{0\leq i\leq n-1}$, and $(w_{i})_{0\leq i\leq n-1}$, such that the following holds: for each $0\leq i\leq n-1$, $v_{i}=n_{11}(u_{i})$, $w_{i}=n_{10}(u_{i})$, and $u_{i+1}=n_{10}(v_{i})=n_{11}(w_{i})$. Similarly, a vertical prop of length $n$ in $G$ is a sequence of distinct vertices $(u_{i})_{0\leq i\leq n}$, $(v_{i})_{0\leq i\leq n-1}$, and $(w_{i})_{0\leq i\leq n-1}$, such that the following holds: for each $0\leq i\leq n-1$, $v_{i}=n_{10}(u_{i})$, $w_{i}=n_{00}(u_{i})$, and $u_{i+1}=n_{00}(v_{i})=n_{10}(w_{i})$ (see Figure 6). The vertices $u_{0}$ and $u_{n}$ are respectively said to be the starting and ending vertex of the (horizontal or vertical) prop.
@@
Let $\mathrm{Pr}_{n}$ denote the graph induced by a (vertical or horizontal) prop of length $n$. That is, $\mathrm{Pr}_{n}$ can be obtained from $n$ disjoint 4-cycles $u_{i}v_{i}u_{i}^{\prime}w_{i}$ ($0\leq i\leq n-1)$ by identifying $u_{i}^{\prime}$ with $u_{i+1}$, for each $0\leq i\leq n-2$. Then $\mathrm{Pr}_{n}$ has $3n+1$ vertices. Consider any fixed embedding of $\mathrm{Pr}_{n}$ in the plane as a unit-square graph. Since $\mathrm{Pr}_{n}$ is triangle-free, this unit-square embedding of $\mathrm{Pr}_{n}$ also gives a planar embedding of $\mathrm{Pr}_{n}$ (with the same circular order of neighbors around each vertex), see Proposition 2.1. There are multiple non-equivalent planar embeddings of $\mathrm{Pr}_{n}$, however a simple area computation shows that in any planar graph drawing coming from a unit-square embedding of $\mathrm{Pr}_{n}$ each 4-cycle of $\mathrm{Pr}_{n}$ is a face, distinct from the outerface, so up to reflection the resulting planar embedding is unique. This implies the following.
@@
Observation 6.4 gives a way to force some squares to align vertically or horizontally. We will now use this to obtain a construction with two specific vertices $v_{1}$ and $v_{2}$, and a sequence of vertices $u_{1},\ldots,u_{n}$ such that in any unit-square embedding, the segment $v_{1}v_{2}]$ is close to a diagonal and $u_{1},\ldots,u_{n}$ are close to be evenly spaced on this segment. We call the latter property almost-perfect rigidity (see the proof below for more details).
@@
  • a cycle $C$ of length $16k+16$, depicted with bold black edges in Figure 7;
@@
  • another copy of $\mathrm{Pr}_{8k+6}$ with endpoints $v_{2},v_{4}\in C$, such that $v_{2}$ and $v_{4}$ are antipodal on $C$ and, for any $1\leq i\leq 4$, $v_{i}$ and $v_{i+1}$ are at distance $4k+4$ on $C$ (where indices are taken modulo 4 plus 1). Moreover the two props intersect in their middle, forming a star on 5 vertices (depicted in bold red edges in Figure 7);
  • on the subpath of $C$ of length $4k+4$ between $v_{1}$ and $v_{2}$, there exists a unique vertex set $I$ of size $k$ in which all vertices are at distance at least 4 from $v_{1}$ and $v_{2}$ on $C$, and any two vertices of $I$ are at distance at least 4 on $C$. For each vertex $v\in I$, we create two copies of a 3 by 3 grid and add an edge between $v$ and a vertex of degree 3 in each of the two grids. The $2k$ copies of the grid are denoted by $H_{1},H_{1}^{\prime},\ldots,H_{k},H_{k}^{\prime}$ in order from $v_{1}$ and $v_{2}$.
@@
By Observation 6.3, $d_{\infty}(v_{1},v_{3})>8k+6$ and $d_{\infty}(v_{2},v_{4})>8k+6$. As $d_{\infty}(v_{i},v_{i+1})\leq 4k+4$ for any $1\leq i\leq 4$, this implies (by the triangle inequality) that $d_{\infty}(v_{i},v_{i+1})>4k+2$ for any such $i$ (with indices taken modulo 4 plus 1). Hence, we have proved that $4k+2<d_{\infty}(v_{1},v_{2})\leq 4k+4$. Using Observation 6.3, this implies that
@@
where $x(\cdot)$ and $y(\cdot)$ denote the $x$- and $y$-coordinates of the points as before. Let us denote by $u_{1},u_{2},\ldots,u_{4k+5}$ the vertices on the subpath $P$ of $C$ of length $4k+4$ between $v_{1}$ and $v_{2}$ (with $u_{1}=v_{1}$ and $u_{4k+5}=v_{2}$). The previous result implies that for any $1\leq i\leq j\leq 4k+5$,
@@
In particular, if we translate and rotate the vertices of the embedding of $G_{k}$ such that $v_{1}$ lies at coordinates $(0,0)$ and $v_{2}$ lies at $\ell_{\infty}$-distance at most 2 from $(4k+4,4k+4)$, then for any $1\leq i\leq 4k+5$, $u_{i}$ lies in the square with corners (i3,i3)(i-3,i-3) and (i1,i1)(i-1,i-1). We call this property the almost-perfect rigidity of $G_{k}$.
@@
The results obtained above show that for any $A,B\subseteq\{1,\ldots,N\}$, $g(L(A),R(B))$ is triangle-free, and it a unit-square graph if and only if $A$ and $B$ are disjoint. As $G_{N}^{\prime}$ and $G_{N}$ have $O(N)$ vertices and the subgraph induced by the closed neighborhood of $S$ is independent of $A$ and $B$ and contains 6 vertices, the class of triangle-free unit-square graphs is $(6,1)$-disjointness-expressing. ∎
@@ 7.3. Stripes
In this section we discuss the properties of triangles and structures called stripes, that will be used later to describe the behavior of unit disks. For $k>0$ and $\delta\in[0,k]$, a triangle in the plane is said to be $(k,\delta)$-almost-equilateral if all sides have length at least $k-\delta$ and at most $k+\delta$. By the law of cosines and the approximation $\arccos(1/2-x)=\pi/3+O(x)$ as $x\to 0$, we have the following.
@@
Assume that $\delta>0$. By Observation 7.1, the maximum angle between the lines $u_{i}v_{i}$ and $u_{i+1}v_{i+1}$ is of order $O(\delta/k)$. Hence, there exists a constant $\alpha$ (independent of $k$ and $\delta$) such that if $\ell=\alpha k/\delta$, the maximum angle between $u_{0}v_{0}$ and $u_{\ell}v_{\ell}$ in any $(k,\delta)$-almost-equilateral embedding is close to $\pi$. In this case, by Lemma 7.2 the minimum (Euclidean) distance $m$ between $u_{0}$ and $v_{\ell}$ in any $(k,\delta)$-almost-equilateral embedding is of order $\Theta(k/\delta\cdot k)=\Theta(k^{2}/\delta)$. Moreover, any $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$ realizing this minimum $m$ is close to some semicircle with endpoints $u_{0}$ and $v_{\ell}$, in the sense that all the vertices of $S_{\ell}$ lie at distance $O(k)$ from the semicircle (see Figure 11). We will need a looser version of this observation in the slightly weaker setting where $d_{2}(u_{0},v_{\ell})\leq m+1$, instead of $d_{2}(u_{0},v_{\ell})=m$.
@@
Let $\delta>0$, $k\gg\delta$ and $\ell=\lceil\alpha k/\delta\rceil$ as above, and let $m=\Theta(k^{2}/\delta)$ be the minimum Euclidean distance between $u_{0}$ and $v_{\ell}$ in any $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$. Consider a $(k,\delta)$-almost-equilateral embedding of $S_{\ell}$ where $d_{2}(u_{0},v_{\ell})\leq m+1$. Let $c$ be the midpoint of the segment $u_{0},v_{\ell}]$. Then no vertex of the stripe $S_{\ell}$ is contained in the disk of center $c$ and radius m/2O(k)m/2-O(k), and all vertices of the stripe $S_{\ell}$ are contained in a disk of center $c$ and radius $O(m)=O(k^{2}/\delta)$.
@@ 7.4. Quasi-rigid graphs
We define $T_{n}$ as follows. We consider two adjacent vertices $u,v$ of the infinite square grid graph and define $X$ as the set of vertices at distance at most $2n+1$ from $u$ or $v$ in the square grid graph, and $Y\subseteq X$ as the set of vertices at distance exactly $2n+1$ from $u$ or $v$. Note that $|X|=2(2n+2)^{2}=8(n+1)^{2}$ and $|Y|=8n+6$. The vertices of $Y$ are denoted by $y_{1},\ldots,y_{8n+6}$. The graph $T_{n}$ is obtained from the subgraph of the square grid graph induced by $X$ by adding, for each vertex $y_{i}$ of $Y\subseteq X$, a vertex $c_{i}$ adjacent to $y_{i}$. We finally add edges to form a cycle $C$ containing all vertices $c_{i}$ in order, together with 10 new vertices in the corners (2 or 3 at each corner, see Figure 12). In Figure 12, the vertices of $C$ are depicted in black, the vertices of $X$ are blue or red, and the vertices of $Y$ are the blue or red vertices that are adjacent to a vertex of $C$.
The cycle $C$ has length $8n+16$. Note that the resulting graph $T_{n}$ contains $8n^{2}+24n+24=O(n^{2})$ vertices and is a planar triangle-free unit-disk graph (see Figure 12). A simple area computation shows that in any unit-disk embedding of $T_{n}$, $C$ bounds the outerface of the corresponding planar graph embedding (by the arguments of Section 2.2 there is a unique planar graph embedding in the sphere, and as all faces except $C$ have size at most 8, $C$ must be the outerface).
@@
Let $f(n)$ be the infimum Euclidean distance between two antipodal vertices $c$ and $c^{\prime}$ of the cycle $C$ in a unit-disk embedding of $T_{n}$. By Lemma 7.4, $f(n)=\Omega(n)$. Since $C$ has length $O(n)$, it follows that $f(n)=\Theta(n)$. Assume for simplicity that the infimum $f(n)$ is a minimum (otherwise we work with a sequence of unit-disk embeddings such that the Euclidean distance between $c$ and $c^{\prime}$ tends to $f(n)$). Let $Z_{n}$ be the point set of a unit-disk embedding of $T_{n}$ in which $d_{2}(c,c^{\prime})=f(n)$. Add $\lceil f(n)\rceil-1$ points, evenly spaced on the line-segment $c,c^{\prime}]$ with respect to the Euclidean distance (note that together with $c$ and $c^{\prime}$, any two consecutive points on $c,c^{\prime}]$ lie at Euclidean distance at most 1 apart). Let $Z_{n}^{\prime}$ denote the resulting point set and $T_{n}^{\prime}$ be the resulting unit-disk graph. Note that $T_{n}^{\prime}$ has $O(n^{2})$ vertices and in any unit-disk graph embedding of $T_{n}^{\prime}$, $f(n)\leq d_{2}(c,c^{\prime})\leq f(n)+1$, with $f(n)=\Omega(n)$.
@@
Assume we have found a family $\mathcal{T}^{1+c}$ which is quasi-rigid with density $g(k)=O(k^{1+c})$, for some $c\geq 0$. Take some graph $G\in\mathcal{T}^{1+c}$ with $g(k)$ vertices and that has two vertices $x$ and $y$ such that in any unit-disk embedding of $G$, $k\leq d_{2}(x,y)\leq k+1$. Let $X$ be the point set associated to some unit-disk embedding of $G$, and let $d:=d_{2}(x,y)$ in this point set ($k\leq d\leq k+1$). Consider a $(d,0)$-almost equilateral embedding of a stripe $S_{\ell}$ (for some integer $\ell\geq 0$ whose value will be fixed later), and for each edge $uv$ in $S_{\ell}$, add an isometric copy of $X$ in which $x$ is identified with $u$ and $y$ is identified with $v$ (by definition of $X$ we can always find such an isometric copy of $X$). See Figure 13 for an illustration. We denote by $X^{(\ell)}$ the resulting point set, and by $G^{(\ell)}$ the resulting unit-disk graph. Note that the different copies of $X$ interact and thus $G^{(\ell)}$ contains more edges than the unions of copies of $G$. Fix any unit-disk embedding of $G^{(\ell)}$, and consider the points corresponding to $x$ and $y$ in each of the copies of $G$ in $G^{(\ell)}$. By the properties of $G$, the union of all these points form a $(d,1)$-almost equilateral embedding of a stripe $S_{\ell}$. Let $a$ and $b$ be the ends of this stripe. By Lemma 7.3, there exists a constant $\alpha>0$ such that if we set $\ell=\lceil\alpha d\rceil=O(k)$, the following holds. Let $m$ be the minimum distance between $a$ and $b$ in any unit-disk embedding of $G^{(\ell)}$. Then $m=\Theta(d^{2})=\Theta(k^{2})$. As before we consider a point set realizing this distance and we add $\lceil m\rceil-1$ evenly spaced points on the segment $[a,b]$. We denote the resulting unit-disk graph by $H$ and we note that in any unit-disk graph embedding of $H$, $m\leq d_{2}(a,b)\leq m+1$. The point set $X^{(\ell)}$ has size at most $5\ell g(k)=O(k^{2+c})$ as there are $4\ell+1$ edges in $S_{\ell}$. We have added at most $m=O(d^{2})=O(k^{2})$ vertices to the graph, so $H$ has $O(k^{2+c})$ vertices. As $m=\Omega(k^{2})$, the family of all graphs $H$ created in this way is quasi-rigid with density $g^{\prime}(k)=O(k^{1+c/2})$.
@@
For any $\varepsilon>0$ there is a family of unit-disk graphs which is quasi-rigid with density $O(k^{1+\varepsilon})$. More precisely, for any sufficiently large $k$ there is a unit-disk graph $G$ with $O(k^{1+\varepsilon})$ vertices with two vertices $a$ and $b$ such that in any unit-disk embedding of $G$, $k\leq d_{2}(a,b)\leq k+1$, and $a$ and $b$ are joined by a path $P$ of length at most $k+1$ such that at least $k/2$ consecutive vertices of $P$ lie at Euclidean distance at least $\Omega(k)$ from GPG-P.